find the least no which when divided by 17 ,16,15 and the remainder is 3.10,0 respectively 

Grade:12

2 Answers

Aman Bansal
12 years ago

Dear Keshav,

x = LCM(15,16,17) - 15 will be divisible by 15
(x - 1) = LCM(15,16,17) - 16 will be divisible by 16
(x - 2) = LCM(15,16,17) - 17 will be divisible by 17

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


keshav kr
12 years ago

plz provide answer that is not ans